OpenStudy (anonymous):

how do you solve cos2x = sinx when the directions tell you to find all solutions between the interval (0,2pi)?

OpenStudy (anonymous):

By rewriting the above equation: \[\cos(2x) = \cos(\frac{ \pi }{ 2 } - x)\] and then solving: \[2x = \pm(\pi/2 -x) +2*\pi*k (k = 0, \pm1, \pm2 ...)\]

OpenStudy (anonymous):

You thus get 2 expressions: \[2x = \pi/2 - x +2*\pi*k \] and \[2x = -\pi/2 + x + 2*\pi*k\] which you need to solve for x and then substitute k =0, 1 -1 etc until you get results higher then 2pi or lower than 0